OpenStudy (doc.brown):

How fast is a particle of 173kg moving if it has 960J of energy?

ganeshie8 (ganeshie8):

use kinetic energy formula

ganeshie8 (ganeshie8):

\(\large KE = \frac{1}{2} mv^2\)

OpenStudy (doc.brown):

\[E_{K}=\frac{1}{2}mv^2\]\[v=\sqrt{\frac{2E_{K}}{m}}\]\[v=\sqrt{\frac{2(960J)}{173kg}}\]\[v=3.3\frac{m}{s}\]\[3.3\frac{m}{s}\times\frac{3.6km\cdot s}{m\cdot h}=12\frac{km}{h}\]Is this right?

ganeshie8 (ganeshie8):

3.33 m/s is correct, after that... if ur arithmetic is correct, it shud be correct :)

ganeshie8 (ganeshie8):

btw, why we're ignoring PE ?

OpenStudy (doc.brown):

I wanted to see how fast a heavier particle would have to go to have the same energy as 12g at 400m/s. Why? What would you do with potential energy?

ganeshie8 (ganeshie8):

okay, total energy = PE+KE

ganeshie8 (ganeshie8):

\(\large \frac{1}{2} m_1v_1^2 = \frac{1}{2} m_2v_2^2\)

ganeshie8 (ganeshie8):

if you knw the masses, you can find the velocity